OpenStudy (anonymous):

Two cards drawn from standard deck of cards. P( 2 aces)

OpenStudy (anonymous):

let me send you the file

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

What you think?

OpenStudy (anonymous):

I am going to say \[\frac{4}{52} \cdot \frac{3}{52}\]

OpenStudy (anonymous):

There are 4 aces in a deck of cards

OpenStudy (anonymous):

doesn't say replacement or no replacement,that why I'm confusing

OpenStudy (anonymous):

You pick one ace and there are 3 left

OpenStudy (anonymous):

Can you do all? ty

OpenStudy (anonymous):

any one ? do as much as you can , medal help

OpenStudy (anonymous):

The deck has 52 cards. The first time it is 4/52 and the second 3/51. So it is 12/2652 or 1/221

OpenStudy (anonymous):

adriano what number?

OpenStudy (anonymous):

told you. 1/221

OpenStudy (anonymous):

not answer,what number you are solve ?

OpenStudy (anonymous):

i dont get you

OpenStudy (anonymous):

I get it now ,TY

OpenStudy (anonymous):

thank your guy I will take a look, medal will send

OpenStudy (anonymous):

4/52*3/51= 12/2652 =1/221 ( in case the card pulled is not returned in cas return allowed it will be 4/52*4/52=16/2704=1/169